Grade 12th passInorganic Chemistry

Which oxoacid of chlorine exibits +4 , +6 oxidation state?

Sayeed Anwar
7 Years agoGrade 12th pass

1 Answer

Subham Patel
7 Years ago
We have 4 oxoacids of chlorine -
1) hypochlorous acid ( HoCl) 
= This have oxidation state of +1 
2) Chlorous acid (HClO2)
         H  - O - Cl=O is the structure 
  Thus oxidation state is = 1-1+x-2=0
                                                X=2
3)chloric acid (HClO3)
     H - O - Cl =O
                  O          ( here two double bond with cl and O present) 
Oxidation state is    1-1+X -2-2=0
                                         X=4 
4) perchloric acid ( Hclo4)
            O
H- O  - Cl =O .  ( Here cl and O form three = bond)
            O 
Oxidation state is 1-1+X -2-2-2=O
                                      X=6
Thus , HClO3 and HclO4  have oxidation state 4 and 6 respectively.... 
                                                       Thank u